can you please help with writing up the calculations with formulas for one- way ANOVA for a null hypothesis of... the output from JASP is provided with ANOVA output and descriptive output

Ho : μ1 = μ2

     Ha : μ1

Descriptives - sbp

sex

Mean

SD

N

female

46.464

11.145

56

male

47.864

11.806

44

ANOVA - sbp

Cases Sum of Squares df Mean Square F p
Note. Type III Sum of Squares
sex 48.250 1.000 48.250 0.369 0.545
Residual 12825.110 98.000 130.868
Answer #1

let, A denotes the factor sex. Now, A has two categories Al kerale, mode sizes Ni & Na respectively. Az Ramale) with TotalAgain, omale a t Eis - G - 1 syn = (11.14s)² + (464642. 2) Ey = 56 [(1 245) ? 46.464}] 3 1 278541402 llly Vmale = {yzj - 52 -TSS Syn - Too = (y + 7 ) - Too = [127854.602+ 100135.142] - (47083 n- with df = 2 12 873.36 21000 -aq SS (Sea) - STU - To - (